### v1

# Adaptive Blind All-in-One Image Restoration

Abstract
Blind all-in-one image restoration models aim to recover a high-quality image from an input degraded with unknown distortions. However, these models require all the possible degradation types to be defined during the training stage while showing limited generalization to unseen degradations, which limits their practical application in complex cases. In this paper, we propose a simple but effective adaptive blind all-in-one restoration (ABAIR) model, which can address multiple degradations, generalizes well to unseen degradations, and efficiently incorporate new degradations by training a small fraction of parameters. First, we train our baseline model on a large dataset of natural images with multiple synthetic degradations, augmented with a segmentation head to estimate per-pixel degradation types, resulting in a powerful backbone able to generalize to a wide range of degradations. Second, we adapt our baseline model to varying image restoration tasks using independent low-rank adapters. Third, we learn to adaptively combine adapters to versatile images via a flexible and lightweight degradation estimator. Our model is both powerful in handling specific distortions and flexible in adapting to complex tasks, it not only outperforms the state-of-the-art by a large margin on five- and three-task IR setups, but also shows improved generalization to unseen degradations and also composite distortions.
https://aba-ir.github.io/
1 Introduction
Figure 1: Our model significantly outperforms state-of-the-art all-in-one image restoration (IR) methods, Restormer [63], PromptIR [40], and DiffUIR [66], across five known IR tasks, three unseen tasks, and three mixed degradation scenarios. The plot is normalized along each axis, with the lowest value positioned on the second circle and the highest value on the outermost circle.
Image restoration (IR) is a fundamental task in computer vision, essential for enhancing visual quality and optimizing the performance of downstream tasks [45, 37]. IR aims to reconstruct high-fidelity images by systematically removing various degradations present in low-quality inputs. These degradations often emerge through a complex interplay of external environmental conditions and camera limitations during the image acquisition process such as adverse weather conditions [28, 48, 49], noise [13], blur [41] and low-light environments [4, 68].
The inherently ill-posed nature of IR presents a significant challenge for conventional approaches limiting their effectiveness [38, 22, 34, 21, 47, 12, 18, 15]. Recent advances in deep learning techniques have led to remarkable progress in IR [27, 63, 40, 41, 11], achieving substantial improvements in reconstruction accuracy. However, these frameworks demonstrate superior performance only in modeling dedicated degradations, as they are typically trained on datasets with individual and specific IR tasks.
To address the limitations of using separate models for each degradation type, recent image restoration (IR) models have adopted an all-in-one approach, i.e., they are designed to handle multiple degradation types within a single model, alleviating the dependency on dedicated models for various IR tasks. While these models have demonstrated success in integrating diverse IR tasks into a unified framework, several challenges persist when applied to real-world IR problems. First, these models assume that the type of distortions are known in advance, allowing them to specifically target and remove them [63]. In practice, however, IR must operate in a blind setting, with no prior knowledge of the degradation present in a given image, making blind IR significantly more challenging. Second, real images frequently contain composite degradations — e.g., a moving vehicle in a low-light scene, or a compressed image captured in hazy weather. Although existing models are designed to handle diverse degradation types, they typically process only a single type of distortion at a time, limiting their effectiveness in practical cases. Third, current all-in-one methods require access to all degradation types during training, which restricts their generalizability to unseen degradations. Adapting these models to new, unobserved degradation types while retaining the all-in-one functionality typically necessitates retraining the entire model with the expanded set of degradations — a process that is both computationally expensive and time-consuming.
Contribution: This paper presents an Adaptive Blind All-in-One Image Restoration (ABAIR) method, designed to bridge the gap between IR techniques and their application in practical complex scenarios. ABAIR effectively addresses multiple and composite degradations with a flexible structure that can be easily updated to include new degradations. Our approach is a simple yet effective scheme that combines three main components. First, we propose a large-scale pre-training with synthetic degradations to obtain a robust weight initialization. We propose a modified CutMix [61] to generate images under multiple distortions with a segmentation head for degradation recognition. Second, we learn disentangled restoration modules by training independent adapters — specifically LoRA [19] — to bridge the gap between synthetic and real-world data. Third, to derive a blind all-in-one IR method, we learn a lightweight image-degradation estimator to select the most suitable combination of adapters based on the input image. With the pre-training on synthetic degradations and the blending operation, our model achieves the ability to handle composite distortions within a single image. Moreover, the adapter-based design enables a flexible model structure that can be easily updated to address new distortions. We can simply train a new adapter module for the added degradation and retrain the lightweight estimator — without losing prior knowledge on other IR tasks. Our approach offers a blind all-in-one model for versatile IR, achieving superior performance on all-in-one restoration benchmarks including three- and five- degradation settings, generalizing to three unseen degradations and handling mixed degradations, as shown in Fig. 1.
2 Related Work
Single Degradation Image Restoration
Most prior work on IR has typically considered removing a single type of degradation in an image to recover its clean counterpart. Typical single degradation IR tasks include denoising [13], deblurring [41], deraining [7], dehazing [59], low-light enhancement [4, 68], etc. Though achieving promising progress on individual tasks, these methods are only capable of dealing with a specific type of distortion, thus limiting their generalizability to a wider range of IR scenarios.
All-in-One Image Restoration
Recently, multi-degradation and all-in-one IR approaches have gained significant attention. Multi-degradation methods [63, 5, 62, 60, 39] proposed unified model architectures that are effective on multiple IR tasks. However, these methods are trained in a way that one set of parameters can handle only one specific type of degradation. Therefore, one still needs to allocate different weights (though with the same architecture) for different degradations. Furthermore, to restore a versatile image, the type of degradation has to be known so that the corresponding parameters can be loaded, this non-blind scheme further hinders the efficiency and effectiveness for real-world applications.
Meanwhile, blind all-in-one IR approaches utilize specialized modules to blindly distinguish degradation types. For instance, AirNet [27] uses a contrastive-based encoder to extract a latent degradation representation from the input images. X-Restormer [8] adds a spatial self-attention module to the transformer block to enhance spatial mapping capabilities. IDR [64] learns degradation-specific priors and incorporates them into restoration, while DiffUIR [66] equips a diffusion model with a condition guidance. Despite the ability to blindly process images, these methods are still incapable of dealing with composite distortions or unseen distortions, due to their focus on dedicated degradations and standard IR benchmarks.
Figure 2: General schema of our proposed method. Our method is divided into three phases. In Phase I we pre-train our baseline model with synthetic degradations of high-fidelity images. Each image contains different degradations in different regions, and a segmentation head learns to predict them, while a restoration loss aims at restoring the image. In this way, the model is able to distinguish and generalize well to multiple degradations. In Phase II, we learn degradation-specific adaptors using standard image restoration datasets. In Phase III, we learn a lightweight degradation estimator to adaptively blend the adapters based on the degradation profile of the input image. This 3-phase methodology makes our method flexible to deal with images containing multiple distortions and easy to update for new ones as it only requires training an adapter for the new distortion and retraining the degradation estimator.
Prompt learning techniques, which capture task-specific context, have shown promise in guiding adaptation for vision tasks [20]. Leveraging this concept, recent methods encode degradation-specific information and guide the restoration model to enhance low-fidelity images with more than one possible distortion. PromptIR [40], for instance, integrates a dedicated prompt block to capture degradation-specific features from input images, while DA-CLIP [30], MPerceiver [2], ProRes [31] and Painter [50] leverage large pre-trained models as prompt generators. The latter approaches, however, are often constrained by high memory demands of large models. Additionally, existing all-in-one approaches require all degradation types to be predefined during training, limiting the addition of new distortion to already trained models.
In contrast, aiming at versatile image restoration, our model is developed to blindly process images with various kinds of distortion, either in single or composite forms. With the mixed design of lightweight degradation adapters, our model is also flexible and extensible to unseen distortions, thus shows great generalizability towards the challenging blind all-in-one IR problem.
(a) Rain
(b) Haze
(c) Noise
(d) Blur
(e) Low-light
Figure 3: Examples of our synthetic degradation generation for five traditional distortions.
Parameter Efficient Fine-Tuning
Fine-tuning models by updating all parameters is computationally inefficient, especially for large-scale models. Parameter Efficient Fine-Tuning (PEFT) addresses this by reducing the number of trainable parameters and memory usage while achieving comparable performance to full fine-tuning [54]. Low-Rank Adaptation (LoRA) [19] introduced a reparametrization strategy for fine-tuned weights, where the weights of a specific layer $W^{\prime}=W+\Delta W$ are represented as a linear combination of the pre-trained weights $W$ and the updated term $\Delta W$ obtained through a low-rank decomposition, $\Delta W=BA$. Here, $B\in\mathbb{R}^{d\times k}$, $A\in\mathbb{R}^{r\times k}$ with rank $r\ll\text{min}(d,k)$. By optimizing only the low-rank matrices $A$ and $B$, LoRA reduces the number of parameters required for adapting that layer (vs tuning all parameters), minimizing memory and computational demands. Building upon LoRA, alternative decomposition methods have been proposed, such as Vector-based Random Matrix Adaptation [24] and Conv-LoRA[67].
PEFT techniques have proven valuable not only for domain adaptation in large models but also for applications like task arithmetic and continual learning, mitigating the problem of catastrophic forgetting—where models lose previously acquired knowledge upon learning new tasks[10]. In IR, Park et al. [39] applied low-rank decomposition within single-task IR frameworks to enhance model performance. In contrast, we propose a simple but novel approach to merge different adaptors — LoRAs — to obtain a flexible blind all-in-one model capable of learning new tasks without forgetting previously learned IR tasks. This is achieved through a disentangled scheme that requires minimal retraining for each new task.
3 Method
Figure 2 presents an overview of our proposed approach for enhancing low-quality images by systematically addressing the key limitations in current IR methods. Our approach consists of three phases, roughly targeting the following challenges: (i) robust generalization to varying types of degradations (Phase I), (ii) effective adaptation to specific degradations (Phase II), (iii) flexible all-in-one mechanism able to handle unknown and mixed degradations (Phase III). Note that we freeze the parameters of baseline after Phase I and adapters after and II, thus preserving maximal knowledge learned from previous tasks and avoid catastrophic forgetting. This design also allows us to adapt to new distortions with minimal additional training as only a new adapter and the light-weighted estimator will be trained.
Phase I: Pre-training with synthetic degradations.
Recent advances demonstrate that large-scale pre-trained models can significantly improve performance across various tasks [1]. Moreover, pre-training with synthetic data has also shown effectiveness when the domain gap is properly addressed [55]. We hypothesize that applying large-scale pre-training to IR can yield a notable performance boost compared to training solely on traditional IR datasets. Thus, as depicted in Fig. 2, Phase I involves training a baseline IR model with large-scale data with synthetic degradations to provide a robust weight initialization for handling complex degradations.
To obtain large-scale data for pre-training, we define a degradation pipeline that generates a low-quality counterpart for each high-quality input image by introducing different synthetic distortions. We focus on five common distortions: noise, blur, rain, haze, and low-light conditions. Each synthetic distortion closely approximates its real-world counterpart. For instance, low-light conditions are simulated by compressing the image histogram, reducing the dynamic range of pixel intensities. In contrast, haze is introduced by adding an achromatic layer based on the depth map estimated from DepthAnythingv2 [55]. Figure 3 illustrates examples of each synthetic degradation. We use the Google Landmarks dataset (GLD) [53], a large-scale collection of five million images of diverse landmarks worldwide. To ensure that the inputs to our pipeline are high quality, we filter out the images with a resolution smaller than 400 pixels on any side and exclude images with a NIMA score [46] below the 4.90 threshold, retaining only high-quality natural images.
To construct our baseline model, we build on recent advancements in all-in-one IR methods. Particularly, we use the Restormer architecture [63] with spatial attention mechanisms [8] and incorporate a modified version of PromptIR [40] prompt blocks. This combination enables the baseline model to effectively capture the inherent information of various degradations, establishing a strong foundation for generalized IR tasks.
While training on single distortions enables the model to address each type individually, real-world IR often involves mixed degradations. However, directly applying multiple distortions to an image often leads to severe quality degradation, making it difficult for the model to learn to reconstruct the original image. To tackle this, we implement two strategies: (i) a degradation CutMix [61] technique, and (ii) a segmentation head with a cross-entropy loss. Our first strategy — inspired by Yang et al. [55] for depth estimation — applies two different degradations divided into different regions. This setup helps the model distinguish and manage multiple degradation types within the same image. As our second strategy, we incorporate a segmentation head into the baseline model, which outputs a per-pixel map of the distortions. By comparing this map with the ground truth, we guide the model to recognize and differentiate multiple types of degradation within a single image. We report additional details about our baseline and the generation of the synthetic distortions in the Supplementary Material.
Table 1: 5-degradations setup. Quantitative results on five IR datasets comparing the state-of-the-art all-in-one methods and our approach. Ours (Oracle) is an upper bound for our approach: it computes the best reachable value in case our estimator always chooses the correct degradation.
PSNR/SSIM Deraining Dehazing Denoising Deblurring Low-Light Average Param.
Rain100L SOTS (Out) BSD68 σ=25 GoPro LoLv1
AirNet [27] 32.98 .951 21.04 .884 30.91 .882 24.35 .781 18.18 .735 25.49 .847 9M
Uformer [51] 35.48 .967 27.20 .958 30.59 .869 26.41 .809 21.40 .808 28.21 .882 52M
IDR [64] 35.63 .965 25.24 .943 31.60 .887 27.87 .846 21.34 .826 28.34 .893 15M
X-Restormer [8] 35.42 .968 27.58 .959 30.92 .880 27.54 .835 20.88 .817 28.47 .891 26M
DA-CLIP [30] 35.49 .970 28.10 .962 30.42 .859 26.50 .807 21.94 .817 28.49 .880 174M
DiffUIR [66] 35.52 .969 28.17 .964 30.92 .879 26.99 .821 20.92 .789 28.50 .880 36M
Restormer [63] 35.56 .970 27.94 .962 30.74 .875 26.84 .818 21.74 .815 28.56 .888 26M
PromptIR [40] 35.40 .967 28.26 .965 30.89 .872 26.55 .808 21.80 .815 28.58 .885 36M
Ours OH 37.73 .978 33.46 .983 31.38 .898 29.00 .878 24.20 .865 31.15 .920 59M
Ours SW 37.79 .979 33.48 .984 31.38 .898 29.00 .878 24.19 .865 31.17 .921 59M
Ours (Oracle) 39.09 0.981 33.54 .984 31.40 0.901 29.10 .879 24.45 .866 31.39 .922 59M
Phase II: Single-task adaptation.
By generating distortions synthetically, we achieve a robust weight initialization for our baseline model. However, a domain gap persists between our pre-training data (due to both the different dataset and the synthetic distions), standard IR datasets, and real-world conditions. For instance, accurately simulating haze is particularly challenging due to its dependence on light scattering by particles, which varies with depth and atmospheric conditions. To bridge this gap, we propose adapting the baseline model with adapters learned for each specific task in its standard dataset, representing Phase II of our approach, as outlined in Fig. 2.
In our approach, each linear and convolutional layer is augmented with a set of LoRA, one per type of degradation, parametrized by $\left\{A_{n}\right\}$ and $\left\{B_{n}\right\}$ where $n$ is the index of the degradation. Following LoRA, these low-rank matrices adjust the initial frozen weights through linear combinations, as described earlier. For simplicity, we omit the subindex indicating the specific layer of the model, but each layer has a separate set of adapters. Due to the proposed pre-training approach and the subsequent specialization via adaptation, our approach achieves superior performance across diverse tasks and generalizes effectively.
Phase III: Multi-task integration.
While LoRA adapters serve as a plug-and-play solution for pre-trained models, there is still the need to select the most suitable adapter based on the input image, when the type of degradation is unknown. To address this limitation and derive a blind all-in-one IR method, we propose to use a lightweight degradation estimator $p\left(n|x;\theta\right)$ to estimate the probability of each degradation $n$ given the input image $x$, parametrized by $\theta$. This estimator is trained on the combination of all (seen) datasets, thus learning to identify the type of degradation present in the image. As shown in Fig. 2, the probability of the estimator is used as weights to linearly combining the task-specific adapters with the baseline parameters. Specifically, given an input image $x$, the weight update for a specific layer with baseline weights $W$ and adapter weights $\left\{A_{n}\right\}$ and $\left\{B_{n}\right\}$ is computed as:
$W^{\prime}\left(x\right)=W+\sum_{n=1}^{N}p\left(n|x;\theta\right)B_{n}A_{n},$ (1)
where $N$ is the number of seen degradations.
We propose two variants of the estimator: (i) one-hot and (ii) soft-weights. The one-hot variant simply selects the adapter corresponding to the degradation with the maximum probability (equivalent to 1 with one-hot probabilities), while the soft-weight variant uses computes the weighted average as in Eq. 1.
Our approach allows the model to linearly combine the degradation-specific knowledge to address both specific and multiple distortions. Its modular and parameter efficient architecture also allows to add additional degradations with small training effort: train additional adapters for the new task, and update the estimator.
Table 2: 3-degradations setup. Quantitative results on three IR datasets comparing the state-of-the-art all-in-one methods and our approach.
PSNR/SSIM Deraining Dehazing Denoising Average
Rain100L SOTS (Out) BSD68 σ=15 BSD68 σ=25 BSD68 σ=50
DL [14] 32.62 .931 26.92 .931 33.05 .914 30.41 .861 26.90 .740 29.98 .875
MPRNet [62] 33.57 .954 25.28 .954 33.54 .927 30.89 .880 27.56 .779 30.17 .899
AirNet [27] 34.90 .967 27.94 .962 33.92 .933 31.26 .888 28.00 .797 31.20 .909
Restormer [63] 35.56 .969 29.92 .970 33.86 .933 31.20 .888 27.90 .794 31.69 .911
PromptIR [40] 36.37 .972 30.58 .974 33.98 .933 31.31 .888 28.06 .799 32.06 .913
Ours OH 38.58 .981 33.71 .985 33.95 .934 31.29 .889 28.04 .798 33.11 .918
Ours SW 38.52 .980 33.62 .984 33.95 .933 31.24 .889 28.01 .796 33.07 .916
Table 3: Quantitative results on additional test datasets with the learned degradations.
PSNR/SSIM Deraining Deblurring Low-Light
Rain100H HIDE Lolv2-Real
IDR [64] 11.32 .397 16.83 .621 17.61 .697
X-Restormer [8] 14.08 .437 25.40 .801 25.42 .876
DiffUIR [66] 14.78 .487 23.98 .739 26.12 .861
Restormer [63] 14.50 .464 24.42 .781 27.12 .877
PromptIR [40] 14.28 .444 24.49 .762 27.70 .870
Ours OH 21.69 .692 27.04 .850 28.09 .907
Ours SW 19.37 .594 27.05 .850 28.09 .906
4 Experiments
We evaluate our method on two setups for all-in-one IR: five- and three-degradation setups. Additionally, we test on datasets excluded from training, novel IR degradation types, and mixed degradation scenarios. The accuracy of the methods is assessed using two well-established metrics: PSNR and SSIM. In all cases we report the mean value for all the test images, and we highlight best and second-best values for each metric. Our approach is compared against recent all-in-one IR methods. Notably, some current methods are either evaluated under different setups or lack available code and models, which complicates direct comparisons. Therefore, we train three state-of-the-art methods, Restormer [63], PromptIR [40], and X-Restormer [8], on the five-degradation IR setup to provide a fair comparison. We used the authors’ provided code. For our approach, we evaluate the two variations based on the variants of the estimator, namely one-hot (OH) and soft weights (SW).
Implementation details:
Our training is conducted in three phases using Adam [23] with weight decay [29], an initial learning rate of $2\times 10^{-4}$, a cosine learning rate scheduler, and a warmup start of one epoch. In phase I, we apply a weighted combination of cross-entropy loss $L_{CE}$ for per-pixel degradation maps and a reconstruction loss $L_{R}$, where $L_{R}$ consists of $L_{1}$ loss and SSIM loss. Both the cross-entropy and SSIM components are weighted with 0.5 In phase II, we use only the reconstruction loss, where, as before, SSIM loss is weighted with 0.5. In phase III, we apply cross-entropy loss between the regressor’s output and the input image’s degradation type.
Table 4: Quantitative results for unseen IR tasks. Note that the models have not been trained for these degradations. Ours* shows results for the lightweight re-training scenario. New adapters are trained for the new tasks and the estimator is retrained with 8 tasks (5-IR case + 3 new ones; only 8M training parameters).
PSNR/SSIM 4-to-8 bits JPEG Q20 Desnowing
Live1 Live1 City-Snow
IDR [64] 24.02 .738 26.51 .913 18.00 .649
X-Restormer [8] 24.73 .745 26.86 .922 18.51 .681
DiffUIR [66] 24.68 .743 26.88 .921 18.39 .671
Restormer [63] 24.64 .743 26.90 .929 18.14 .655
PromptIR [40] 24.70 .740 26.60 .920 18.49 .673
Ours OH 25.25 .742 29.20 .931 18.71 .684
Ours SW 25.32 .743 29.35 .926 18.67 .683
Ours OH 29.14 .826 30.82 .943 24.19 .797
Ours SW 29.03 .810 30.71 .939 24.02 .779
5-Degradation blind IR:
For the blind five-task setup, we follow the protocol by Zhang et al. [64]. Specifically, we use Rain200L[56] for deraining, RESIDE [25] for dehazing, BSD400 [33] and WED [32] for denoising with $\sigma=25$, GoPro [35] for deblurring, and LOL [52] for low-light enhancement. For evaluation, we employ Rain100L [56], SOTS-Outdoor [25], BSD68 [33], GoPro [35], and LOL [52]. Results are reported in Tab. 1. Our approach outperforms all state-of-the-art methods across all tasks, with the exception of PSNR for denoising. Notably, our method achieves a substantial improvement over the state of the art, yielding a 2.91 dB average PSNR gain across the five tasks. In particular, our method surpasses IDR [64] by 2.16 dB on image deraining and provides a 2.40 dB improvement over PromptIR [40] on low-light image enhancement.
GoPro [35] Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
LoLv1 [52] Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Rain100H [56] Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input Restormer [63] PromptIR [40] Ours-OH Ground truth
Figure 4: Qualitative results for single degradation removal, including deblurring on the GoPro [35] dataset, denoising on the LoLv1 [52] dataset, and deraining on the Rain100H [56] dataset.
3-Degradation Blind IR:
Following Li et al. [27], we further evaluate our approach on a 3-task blind IR setup, comparing our approach with specialized all-in-one methods for deraining, dehazing, and denoising. Compared to the five-task setup, we omit deblurring and low-light enhancement and instead introduce two additional noise levels: $\sigma=15$ and $\sigma=50$. Two LoRA adapters for these noise levels are trained, and the estimator is retrained for the new setting. Results for this 3-task setup are presented in Tab. 2. Our approach outperforms all the other state-of-the-art methods on average and stays consistent in all the degradation types.
JPEG [42] Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
4-to-8 bits [42] Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input PromptIR [40] Ours-SW Ours-SW retrained Ground truth
Figure 5: Qualitative results for unseen IR tasks, including JPEG artifact removal and 4-to-8 bit reconstruction. PromptIR [40] and Ours are not trained for this task, while Ours retrained has a specified LoRA in an 8-degradation setup.
Additional test sets:
To assess the generalization capabilities of our approach, we evaluate the model on three datasets not used during training. Specifically, we use the Rain100H [56] dataset for heavy deraining, HIDE [43] for human-centric deblurring, and LoLv2-Real [58] for low-light image enhancement. The results of the three best methods on the 5-degradation setup and our approach are presented in Tab. 3. Our method outperforms all other approaches across all datasets and metrics. Notably, we achieve a significant improvement of 7.19 dB over Restormer [63] on Rain100H [56]. This performance boost can be attributed to the pre-training on synthetic data, which allows our model to learn from a broader set of natural images and degradation scenarios compared to the other methods’ small training set, like Rain200L [56]. Additionally, we achieve over 1 dB PSNR improvement on both HIDE and LoLv2-Real, further demonstrating the effectiveness and robustness of our approach.
Unseen IR tasks:
Although our approach was initially trained on five types of degradations, we further evaluate its generalization to three additional IR tasks not encountered during training. Specifically, we assess our method on JPEG artifact removal and bit-depth reconstruction using the Live1 dataset [42], and on desnowing using CityScapes-Snow-Medium [65]. Results for the top four previous state-of-the-art methods from the five-degradation setup, alongside our approach, are presented in Tab. 4. Remarkably, our method outperforms all other approaches across all tasks and metrics. To show the adaptability of our approach, we additionally train a separate adapter for each new task, along with a new estimator that blends all eight task-specific adapters, thereby expanding the model’s versatility. The last two rows of Tab. 4 report results for the newly integrated tasks — shown with a . Notably, with the addition of these task-specific adapters, our model achieves a notable boost in performance, while maintaining an average PSNR of 30.08 dB across the original five IR tasks, outperforming the other state-of-the-art methods trained on the initial setup.
Blur&Noise Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Haze&Snow Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input Restormer [63] PromptIR [40] Ours-SW Ground truth
Figure 6: Qualitative results for two examples of mixed degradations. The first row depicts an image with blur and noise from the GoPRO dataset [35], while the second row shows an image with haze and snow from the SRRS dataset [6]. Columns display the input image, results from Restormer [63], PromptIR [40], our method, and the ground truth, respectively.
Mixed degradations:
The limited availability of datasets with mixed degradations presents a challenge for training IR models under these complex conditions. To evaluate our approach, we evaluate it on three mixed-degradation scenarios: blur combined with noise $\sigma{=}25$ using the GoPro dataset[35], blur with JPEG artifacts using the REDS dataset [36], and haze combined with snow using the SRRS dataset [6]. Notably, some degradations, such as JPEG artifacts and snow, are unseen during training, further increasing the difficulty of these tests. As shown in Tab. 5, our model consistently outperforms other methods across all scenarios and metrics. Handling blur with noise is particularly challenging due to their contrasting nature: blur affects low frequencies, while noise affects high frequencies. Even though, our model achieved around 2dB improvement while others remained at the level of 22dB.
Table 5: Quantitative results on datasets with mixed degradations.
PSNR/SSIM Blur&Noise Blur&JPEG Haze&Snow
GoPro REDS SRRS
IDR [64] 21.98 .683 23.02 .681 20.51 .789
X-Restormer [8] 22.67 .669 23.98 .710 20.76 .805
DiffUIR [66] 22.71 .670 24.00 .711 20.86 .802
Restormer [63] 22.35 .662 23.24 .698 20.76 .800
PromptIR [40] 22.89 .671 23.92 .705 20.94 .803
X-Restormer [8] 22.67 .669 23.98 .710 20.76 .805
Ours OH 24.30 .743 24.81 .717 21.48 .834
Ours SW 25.14 .750 24.97 .719 22.09 .839
Table 6: Ablation studies on types of pre-training for Phase I, and the rank of LoRA [19] for Phase II and III.
Pre-training PSNR SSIM
IR datasets 28.50 .892
GLD+synth. 30.63 .913
+ CutMix 31.09 .920
+ Aux. segm. 31.17 .921
Rank PSNR SSIM Params
4 31.17 .921 3.6M
8 31.14 .920 7.2M
16 30.97 .916 14.3M
Qualitative results:
In Fig. 4, we present examples featuring three types of single degradations from the five-task setup on which the model was trained. Our approach effectively enhances the license plate in the first row, reduces noise in low-light conditions in the second row, and removes heavy rain streaks in the third. In Fig. 5, we show unseen IR tasks including JPEG artifact removal and 4-to-8-bit reconstruction using the Live1 dataset [42]. Our method successfully removes JPEG artifacts from the parrot’s plumage and artifacts caused by bit-depth reduction from the cloud and blue sky. By training a new adapter for these tasks (ours-SW retrained), we achieve superior results with minimal additional training time. Finally, in Fig. 6, we demonstrate our method’s performance under mixed degradations. The first row shows that our method achieves the best reconstruction of the text in the image, while the second row illustrates our model’s effectiveness in removing both haze and snowflakes.
Ablation studies:
In Tab. 6, we perform an ablation study to assess the contribution to the final performance of the proposed pre-training pipeline and the LoRAs’ rank. For pre-training in Phase I, we use the same setup as the five-task configuration, i.e. training the LoRAs and the estimator using task-specific IR datasets and the obtained pre-trained weights. The first row — IR datasets — reflects pre-training on standard IR benchmarks, yielding performance comparable to state-of-the-art methods — see Tab. 1. In contrast, simply pre-training on GLD with synthetic degradations improves performance by 2.15 dB PSNR, demonstrating the importance of Phase I to provide the model with strong generalization capabilities. Further incorporating both the CutMix-like [61] strategy and the auxiliar segmentation task provides an additional performance boost. We also compare LoRA [19] with different ranks for Phase II and III. The model achieves the best performance and efficiency when the rank is set to 4, which we adopted as our default setting. For more ablation results, please refer to our Supplementary Material.
5 Conclusion
In this work, we introduced an adaptive blind all-in-one IR model aiming towards practical IR. We design both dedicated adapters for strongly handling specific distortions but also a flexible architecture for dealing with practical IR challenges.We first developed a pre-training pipeline featuring multiple synthetic degradations over a large dataset that boosts the generalization of the model. Second, we derived compact per-task adapters that robustly adapt to specific degradations. Third, we developed a lightweight degradation estimator that identifies varying degradations to blend the respective adapters, which is also capable of efficiently incorporate new degradations by training a small fraction of parameters. Our model largely outperforms the state-of-the-art on five- and three-task IR setups, and shows improved generalization to unseen datasets and IR tasks.
Acknowledgements
DSL, LH, and JVC were supported by Grant PID2021-128178OB-I00 funded by MCIN/AEI/10.13039/ 501100011033 and by ERDF ”A way of making Europe”, by the Departament de Recerca i Universitats from Generalitat de Catalunya with reference 2021SGR01499, and by the Generalitat de Catalunya CERCA Program. DSL also acknowledges the FPI grant from Spanish Ministry of Science and Innovation (PRE2022-101525). LH was also supported by the Ramon y Cajal grant RYC2019-027020-I. SS was supported by the HORIZON MSCA Postdoctoral Fellowships funded by the European Union (project number 101152858)
6 Supplementary Material
We provide additional material to supplement our main submission. Specifically, we cover:
  1. A.
    Details on our baseline architecture.
  2. B.
    Description of the generation process for synthetic degradations, including rain, haze, noise, blur, and low-light conditions. Additional dataset examples.
  3. C.
    Additional ablation experiments on LoRA’s rank and other decomposition methods.
  4. D.
    Details on our estimator architecture and additional analysis of the estimator’s performance.
  5. E.
    Additional qualitative results on known datasets, unseen tasks, and images with mixed degradations.
Appendix A Baseline Architecture Details
We present an Adaptive Blind All-in-One Image Restoration (ABAIR) method, designed to bridge the gap between IR techniques and their application in practical complex scenarios. Our approach follows a three-phase scheme. The first phase involves pre-training an IR baseline using natural images with synthetic degradations. In this section, we describe this baseline architecture.  Figure 7 shows the details of our baseline model. Our baseline model adopts the Restormer [63] architecture, a transformer-based UNet-like framework. Given a degraded image, the model first applies a convolutional layer to extract low-level features of size $H\times W\times C$, where $H$ and $W$ are the spatial dimensions, and $C{=}48$ in all our experiments. These features are then processed through a four-level encoder-decoder structure composed of transformer blocks, with pixel unshuffling and shuffling [44] used for downsampling and upsampling, respectively. Finally, a convolutional layer generates the residual image, which is added to the degraded input image to produce the restored output.
Each transformer block comprises a channel self-attention module followed by a spatial self-attention module. For channel self-attention, we adopt the implementation from Zamir et al. [63], while for spatial attention, we use the overlapping cross-attention mechanism proposed by Chen et al. [9]. This combination effectively addresses the limitations of Restormer’s U-shaped architecture, particularly its difficulty in reconstructing high-frequency details [8]. Additionally, inspired by Potlapalli et al. [40], we integrate a Prompt Block between the transformer blocks in the upsampling path. This block facilitates the architecture in identifying degradation-specific features in the input image by combining extracted features with a set of model parameters. The design of our Prompt Block is depicted in Fig. 8.
Figure 7: Overview of our baseline model. The input degraded image is processed through a four-level U-shaped network with transformer blocks. In the upsampling path, prompt blocks are integrated to assist the model in capturing degradation-specific information.
Figure 8: Overview of our Prompt Block. The input features, derived from the output of a transformer block, are modulated through a weighted element-wise multiplication with a set of model parameters. These modulated features are then further refined using an additional transformer block, producing the enhanced output features of the Prompt Block.
Appendix B Synthetic Degradations Generation
The first phase of our approach involves pre-training the baseline model using natural images with synthetically generated degradations, including rain, haze, noise, blur, and low-light conditions. Unlike standard IR datasets, our pipeline introduces degradations dynamically to clean input images, offering greater flexibility by enabling the model to be trained on the same image with diverse degradation types and varying severity levels. Moreover, each degradation type is parametrized by a set of values that constrain the severity of the distortion. During training, these parameters are randomly selected for each forward pass. Specifically, we utilize 450K images from the Google Landmarks dataset [53] — those with NIMA score higher than 4.90 and with short-edge resolution larger than 400, providing a diverse set of real scenes. Figure 9 show examples for the different degradations under different parameter settings. In what follows, we detail the generation process for each type of synthetic degradation.
Rain:
Capturing paired degraded and clean images with rain is inherently challenging, as environmental conditions often vary when capturing the same scene under rainy and clear weather. Consequently, deraining datasets typically simulate rain by creating a set of predefined rain stroke masks, which are then added to input images. However, these datasets often include a limited number of masks [56], leading to potential overfitting on the specific patterns present in the dataset. In our case, aiming to derive a more general set of masks, we consider five adjustable parameters: density, length, angle, drop size, and blending weight.
The process for generating a rainy image $I_{rain}$ begins by creating an empty mask $M$ of the same size as the input image $I_{input}$, which will hold the rain-stroke patterns. The density parameter $d$ determines the number of raindrops, computed as a fraction of the total image pixels. Random coordinates for the starting positions of the drops are generated, ensuring they stay within bounds to accommodate the specified drop size $s$. These coordinates are then used to populate $M$ with raindrops. Next, a motion blur kernel $K$ is constructed to simulate the appearance of rain streaks, based on the specified rain length $l$ and angle $\theta$. The mask $M$ is convolved with $K$, creating streaks that mimic natural rain patterns. Finally, the rain streaks are normalized and expanded into three RGB channels to match the input image dimensions. These streaks are blended with the original image $I_{input}$ using a weight $w$, resulting in the final rain-augmented image $I_{\text{rain}}$. All the input parameters are randomly selected for each forward pass. Mathematically —omitting normalizations for clarity,
$I_{rain}=wI_{input}+(1-w)(K*M).$ (2)
The value ranges are rain density $d$:$[0.005,0.02]$, rain length $l$: $[25,35]$, rain angle $\theta$: $[70,110]$, raindrop size $s$: $[1,3]$, and mask weight $[0.75,1]$.
$d{=}0.005$, $l{=}25$, $\theta{=}70^{\circ}$, min${=}0.2$, max${=}0.7$ $s{=}1$, and $w{=}0.75$ $A{=}140$ $\sigma{=}15$ $d{=}9$ and $\Theta{=}0^{\circ}$ $c{=}0.25$ and $\sigma{=}0.5$
$d{=}0.01$, $l{=}30$, $\theta{=}90^{\circ}$, min${=}0.3$, max${=}0.8$ $s{=}2$, and $w{=}1$ $A{=}160$ $\sigma{=}25$ $d{=}25$ and $\Theta{=}45^{\circ}$ $c{=}0.5$ and $\sigma{=}1$
$d{=}0.01$, $l{=}30$, $\theta{=}90^{\circ}$, min${=}0.3$, max${=}0.8$ $s{=}3$, and $w{=}1$ $A{=}180$ $\sigma{=}25$ $d{=}25$ and $\Theta{=}90^{\circ}$ $c{=}0.5$ and $\sigma{=}1.5$
$d{=}0.02$, $l{=}35$, $\theta{=}110^{\circ}$, min${=}0.4$, max${=}0.9$ $s{=}1$, and $w{=}1$ $A{=}200$ $\sigma{=}50$ $d{=}35$ and $\Theta{=}135^{\circ}$ $c{=}0.25$ and $\sigma{=}1.5$
Rain Haze Noise Blur Low-Light
Figure 9: Additional examples showing our synthetic degradation generation for rain, haze, noise, blur, and low-light conditions. On top of each image, we report the specific parameters used to produce the corresponding degraded output.
Haze:
As with rainy images, capturing paired degraded and clean images under hazy conditions is nearly impossible due to the variability of environmental factors. Inspired by prior works [26, 3, 16], we model haze degradation using the Kochsmieder model [17], which describes how the visibility of distant objects diminishes, vanishing into the horizon as a function of their distance from the observer. We can formulate it as:
$I_{haze}=T\cdot I_{input}+(1-T)\cdot A,$ (3)
where $I_{input}$ is the input clean image, $T$ is the transmission map derived from the estimated depth map, $A$ is the airlight, and $I_{haze}$ is the resulting image with synthetic haze.
We estimate the depth map of the input image using DepthAnythingv2 [55]. This depth map is normalized within the predefined minimum and maximum haze values to generate the transmission map $T$. The depth map is then replicated for each color channel to compute Eq. 3. To ensure variability, all parameters are randomly sampled for each forward pass. The parameter ranges are as follows: minimum haze $[0.2,0.4]$, and maximum haze $[0.7,0.9]$, airlight is unique for the three color channels with values ranging $[140,200]$.
Blur:
Blur in images can arise from various sources, such as motion blur, out-of-focus blur, and lens blur, among others. In this work, we focus on simulating the motion blur due to its relevance in practical applications. The blur effect is introduced by convolving the input image $I_{input}$ with a parametrized kernel $K$, designed to model motion blur along a specific direction. The kernel is defined by its size in pixels ($d$) and its angle ($\Theta$), to simulate the directional streaks characteristics of motion blur. Mathematically,
$I_{blur}=K_{\Theta,d}*I_{input}.$ (4)
The kernel size ($d$) is selected from odd values within the range $[9,35]$, while the angle ($\Theta$) specifies the orientation of the blur in degrees, ranging from $[0,360]$.
Noise:
We use the standard Additive White Gaussian Noise (AWGN) approach for noise. In short, we add to the original image a second image that follows a Gaussian distribution with mean 00 and variance $\sigma$. Mathematically,
$I_{noise}=I_{input}+\mathcal{N}(0,\sigma).$ (5)
Low-Light:
When an image is captured under low-light conditions, cameras amplify the sensor’s signal to capture details, amplifying noise, and leading to grainy or speckled artifacts. Furthermore, the reduced dynamic range limits the ability to capture the range of intensities, resulting in color inconsistencies, and loss of detail in shadows and highlights. To simulate these conditions, we compress the input image histogram by a factor $c$ and add a noise with a meager $\sigma$ value. This process can be expressed mathematically as:
$I_{lol}=I_{input}\cdot c+\mathcal{N}(0,\sigma).$ (6)
The compression factor $c$ ranges in the interval $[0.25,0.5]$ and the $\sigma$ takes values in the interval $[0.5,1.5]$.
Table 7: Ablation study on different low-rank adapters and their rank. Results are the mean for all images. LoRA outperforms both VeRA and Conv-LORA. Lower ranks perform better.
PSNR/SSIM Deraining Dehazing Denoising Deblurring Low-Light Average Adapter Param.
Method Rank Rain100L SOTS (Out) BSD68 σ=25 GoPro LoLv1
LoRA [19] 4 37.79 .979 33.48 .984 31.38 .898 29.00 .878 24.19 .865 31.17 .921 3.6M
8 37.75 .978 33.4 .982 31.39 .898 29.02 .878 24.18 .865 31.15 .920 7.2M
16 37.61 .972 33.21 .977 31.31 .896 28.77 .875 23.96 .862 30.97 .916 14.3M
VeRA [24] 4 37.02 .971 32.67 .972 31.32 .896 28.61 .872 23.78 .580 30.68 .858 460K
8 37.09 .971 32.69 .972 31.32 .896 28.64 .873 23.79 .580 30.71 .858 468K
16 37.04 .970 32.62 .970 31.33 .896 28.62 .872 23.84 .581 30.69 .858 476K
Conv-LoRA [67] 4 37.00 .969 32.55 .971 31.32 .896 28.54 .870 23.70 .576 30.62 .856 3.9M
8 36.94 .968 32.44 .968 31.30 .895 28.48 .868 23.62 .575 30.56 .855 7.5M
Appendix C Additional Phase II Ablation Studies
The second phase of our approach involves training a set of adapters — LoRA [19] in the main manuscript — for each type of degradation. In this section, we extend our analysis to include other low-rank decompositions and their rank. Specifically, we evaluate VeRA [24] and Conv-LoRA [67]. The results for the five-degradation setup across the three adapter types are presented in Tab. 7, which also includes the number of trainable parameters for each adapter in the last column. For a fair comparison, we utilize the same baseline model weights and estimator, while training only the task-specific adapters. Among the methods, LoRA [19] achieves the best overall performance. However, VeRA [24] provides competitive results with significantly fewer parameters, as it estimates only two vectors per layer instead of two low-rank matrices. On the other hand, Conv-LoRA [67] performs worse despite having more parameters, owing to its Mixture-of-Experts approach with convolution layers in the decomposed space. Regarding the decomposition rank, we find that ranks of 4 and 8 consistently outperform a rank of 16 in terms of both accuracy and parameter efficiency.
Table 8: Ablation study on different methods for blending the five degradations task-specific LoRA [19] adapters.
PSNR/SSIM Deraining Dehazing Denoising Deblurring Low-Light Average
Method Rain100L SOTS (Out) BSD68 σ=25 GoPro LoLv1
Sum 19.50 .755 18.80 .745 18.10 .730 18.25 .728 17.70 .715 18.67 .736
Average 30.54 .939 20.87 .855 28.98 .785 21.34 .792 15.49 .673 23.84 .809
Ours OH 37.73 .978 33.46 .983 31.38 .898 29.00 .878 24.20 .865 31.15 .920
Ours SW 37.79 .979 33.48 .984 31.38 .898 29.00 .878 24.19 .865 31.17 .921
Appendix D Additional Phase III Analysis
In the third phase of our method, we train an estimator to select or blend the most suitable set of adapters based on the input image. The estimator architecture comprises four blocks of Conv2D layers, each followed by batch normalization, ReLU activation, and max pooling, culminating with a global average pooling layer and a linear projection. The total parameter count for the estimator is 538K.
We present a confusion matrix in Fig. 10, showing the One-Hot predictions for known degradations —both seen and unseen datasets. The estimator predicts known degradations with probabilities exceeding 90% in most cases. However, for one of the unseen datasets, the Rain100H dataset [56], the prediction probability falls below 90%, likely due to the severity of the rain streaks and their resemblance to haze-related degradations. Notably, even when the estimator selects an incorrect adapter, no significant artifacts or undesired effects are introduced, owing to the robustness of the large-scale pre-training.
Finally, to highlight the importance of our estimator, we perform a last ablation comparing our approach to just adding or averaging the five degradation task-specific adapters. Tab. 8 shows the results of this ablation. When all adapters are added, the $\Delta W$ values become excessively large for the baseline model to effectively handle them. On the other hand, averaging the adapters results in suboptimal restored images, as the model fails to specialize in any particular degradation type.
Figure 10: Confusion matrix for our estimator in test images. First five columns: known degradations with the datasets considered at training. Last three columns: unknown datasets for rain, blur, and low light. Specifically, from left-to-right columns, the datasets are: Rain100L [56], BSD68 [33], SOTS (outdoors) [26], GoPro [35], LoLv1 [52], Rain100H [56], HIDE [43], and LoLv2-Real [57].
Appendix E Additional Qualitative Results
Figure 11 presents additional qualitative results for known degradations, including examples from both seen and unseen datasets. Figure 12 highlights results for unknown degradations, with the corresponding mean absolute error (MAE) map displayed below each image to emphasize the differences. Finally, Figure 13 shows qualitative results for mixed degradation scenarios.
Rain [56] Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Haze [26] Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Noise [35] Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Blur [35] Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Low-Light [52] Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Rain100H [56] Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
LoLv2-Real [58] Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input Restormer [63] PromptIR [40] Ours-OH Ground truth
Figure 11: Qualitative results for known degradation removal, both seen datasets —Rows 1 to 5— and unseen datasets —Rows 6 to 8.
JPEG [42] Refer to caption
Snow [65] Refer to caption
Input/Ground truth Restormer [63] PromptIR [40] Ours-SW Ours-SW retrained
Figure 12: Qualitative results on JPEG artifact removal on Live1 dataset [42] and desnowing on CityScapes-Snow-Medium [65]. We show the mean absolute error map below each image.
Blur&JPEG [36] Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Haze&Snow [6] Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input Restormer [63] PromptIR [40] Ours-OH Ground truth
Figure 13: Qualitative results on mixed degradation scenarios. Specifically, blur and JPEG on the REDS [36] dataset and haze and snow in the SRRS [6] dataset.

### v2

@@
 However, these models require all the possible degradation types to be defined during the training stage while showing limited generalization to unseen degradations, which limits their practical application in complex cases.
-In this paper, we propose a simple but effective adaptive blind all-in-one restoration (ABAIR) model, which can address multiple degradations, generalizes well to unseen degradations, and efficiently incorporate new degradations by training a small fraction of parameters.
-First, we train our baseline model on a large dataset of natural images with multiple synthetic degradations, augmented with a segmentation head to estimate per-pixel degradation types, resulting in a powerful backbone able to generalize to a wide range of degradations.
-Second, we adapt our baseline model to varying image restoration tasks using independent low-rank adapters.
+In this paper, we introduce ABAIR, a simple yet effective adaptive blind all-in-one restoration model that not only handles multiple degradations and generalizes well to unseen distortions but also efficiently integrates new degradations by training only a small subset of parameters.
+We first train our baseline model on a large dataset of natural images with multiple synthetic degradations.
+To enhance its ability to recognize distortions, we incorporate a segmentation head that estimates per-pixel degradation types.
+Second, we adapt our initial model to varying image restoration tasks using independent low-rank adapters.
 Third, we learn to adaptively combine adapters to versatile images via a flexible and lightweight degradation estimator.
-Our model is both powerful in handling specific distortions and flexible in adapting to complex tasks, it not only outperforms the state-of-the-art by a large margin on five- and three-task IR setups, but also shows improved generalization to unseen degradations and also composite distortions.
+This specialize-then-merge approach is both powerful in addressing specific distortions and flexible in adapting to complex tasks.
+Moreover, our model not only surpasses state-of-the-art performance on five- and three-task IR setups but also demonstrates superior generalization to unseen degradations and composite distortions.
 https://aba-ir.github.io/
 1 Introduction
-Figure 1: Our model significantly outperforms state-of-the-art all-in-one image restoration (IR) methods, Restormer [63], PromptIR [40], and DiffUIR [66], across five known IR tasks, three unseen tasks, and three mixed degradation scenarios.
+Figure 1: Our model significantly outperforms state-of-the-art all-in-one image restoration (IR) methods, DiffUIR [81], X-Restormer [8], and AdaIR [10], across five known IR tasks, three unseen tasks, and three mixed degradation scenarios.
 The plot is normalized along each axis, with the lowest value positioned on the second circle and the highest value on the outermost circle.
@@
 Recent advances in deep learning techniques have led to remarkable progress in IR [32, 76, 49, 51, 12], achieving substantial improvements in reconstruction accuracy.
-However, these frameworks demonstrate superior performance only in modeling dedicated degradations, as they are typically trained on datasets with individual and specific IR tasks.
-To address the limitations of using separate models for each degradation type, recent image restoration (IR) models have adopted an all-in-one approach, i.e., they are designed to handle multiple degradation types within a single model, alleviating the dependency on dedicated models for various IR tasks.
+However, these frameworks demonstrate superior performance only in modeling dedicated degradations, as they are typically trained on IR datasets with specific degradations.
+To address the limitations of using separate models for each degradation type, recent IR models have adopted an all-in-one approach, i.e., they are designed to handle multiple degradation types within a single model, alleviating the dependency on dedicated models for various IR tasks.
 While these models have demonstrated success in integrating diverse IR tasks into a unified framework, several challenges persist when applied to real-world IR problems.
-First, these models assume that the type of distortions are known in advance, allowing them to specifically target and remove them [63].
+First, these models assume that the type of degradations are specified in advance, allowing them to target and remove them specifically.
 In practice, however, IR must operate in a blind setting, with no prior knowledge of the degradation present in a given image, making blind IR significantly more challenging.
@@
 Our approach is a simple yet effective scheme that combines three main components.
-First, we propose a large-scale pre-training with synthetic degradations to obtain a robust weight initialization.
-We propose a modified CutMix [61] to generate images under multiple distortions with a segmentation head for degradation recognition.
-Second, we learn disentangled restoration modules by training independent adapters — specifically LoRA [19] — to bridge the gap between synthetic and real-world data.
-Third, to derive a blind all-in-one IR method, we learn a lightweight image-degradation estimator to select the most suitable combination of adapters based on the input image.
-With the pre-training on synthetic degradations and the blending operation, our model achieves the ability to handle composite distortions within a single image.
-Moreover, the adapter-based design enables a flexible model structure that can be easily updated to address new distortions.
-We can simply train a new adapter module for the added degradation and retrain the lightweight estimator — without losing prior knowledge on other IR tasks.
-Our approach offers a blind all-in-one model for versatile IR, achieving superior performance on all-in-one restoration benchmarks including three- and five- degradation settings, generalizing to three unseen degradations and handling mixed degradations, as shown in Fig.
-1.
+First, we propose a large-scale pretraining with synthetic degradations to obtain a robust weight initialization.
+To increase variability, we introduce Degradation CutMix [74], a technique for generating images with controlled mixtures of degradations.
+Additionally, we enhance this pretraining by incorporating a segmentation head to recognize per-pixel degradation types.
+Second, we train independent restoration adapters — specifically LoRA [23] — to bridge the gap between synthetic and real data.
+Third, we train a lightweight image-degradation estimator to select the most suitable combination of adapters, enabling a blind all-in-one IR method.
+By leveraging the synthetic pretraining and the specialize-then-merge scheme, our model achieves handling composite distortions within a single image.
+Moreover, our pretraining and adapter-based design provides a flexible and extensible model structure that can be easily updated to address new distortions.
+By training a new adapter for the added degradation and retraining the lightweight estimator, our approach preserves prior knowledge of other IR tasks without requiring full model retraining.
+This enables a blind all-in-one model for versatile IR, achieving state-of-the-art performance on all-in-one IR benchmarks, and generalizing to unseen and mixed degradations, as shown in Figure 1.
 2 Related Work
@@
 Recently, multi-degradation and all-in-one IR approaches have gained significant attention.
-Multi-degradation methods [63, 5, 62, 60, 39] proposed unified model architectures that are effective on multiple IR tasks.
-However, these methods are trained in a way that one set of parameters can handle only one specific type of degradation.
-Therefore, one still needs to allocate different weights (though with the same architecture) for different degradations.
+Multi-degradation methods [76, 4, 75, 73, 48, 58, 19, 38, 18, 9] proposed model architectures that are effective on distinct IR tasks.
+However, these methods are trained so that one set of parameters can handle only one specific type of degradation.
+Therefore, one still needs to allocate different weights for different degradations.
 Furthermore, to restore a versatile image, the type of degradation has to be known so that the corresponding parameters can be loaded, this non-blind scheme further hinders the efficiency and effectiveness for real-world applications.
-Meanwhile, blind all-in-one IR approaches utilize specialized modules to blindly distinguish degradation types.
-For instance, AirNet [27] uses a contrastive-based encoder to extract a latent degradation representation from the input images.
+Meanwhile, blind all-in-one IR approaches utilize specialized modules to distinguish degradation types blindly.
+For instance, AirNet [32] and OneRestore [20] use a contrastive loss to extract a latent degradation representation from the input images.
 X-Restormer [8] adds a spatial self-attention module to the transformer block to enhance spatial mapping capabilities.
-IDR [64] learns degradation-specific priors and incorporates them into restoration, while DiffUIR [66] equips a diffusion model with a condition guidance.
+AdaIR [10] operates in the frequency domain to distinguish the distinct degradation types while IDR [78] learns degradation-specific priors and incorporates them into restoration.
 Despite the ability to blindly process images, these methods are still incapable of dealing with composite distortions or unseen distortions, due to their focus on dedicated degradations and standard IR benchmarks.
@@
 Our method is divided into three phases.
-In Phase I we pre-train our baseline model with synthetic degradations of high-fidelity images.
-Each image contains different degradations in different regions, and a segmentation head learns to predict them, while a restoration loss aims at restoring the image.
-In this way, the model is able to distinguish and generalize well to multiple degradations.
+In Phase I we pretrain our baseline with synthetic degradations on high-fidelity images.
+Each image contains different degradations in different regions — Degradation CutMix [74], and a segmentation head learns to predict them, while a restoration loss aims at restoring the image.
+In this way, the model can distinguish and generalize well to multiple degradations in a controlled manner.
 In Phase II, we learn degradation-specific adaptors using standard image restoration datasets.
 In Phase III, we learn a lightweight degradation estimator to adaptively blend the adapters based on the degradation profile of the input image.
-This 3-phase methodology makes our method flexible to deal with images containing multiple distortions and easy to update for new ones as it only requires training an adapter for the new distortion and retraining the degradation estimator.
+This three-phase methodology makes our method flexible to deal with images containing multiple degradations and easy to update for new ones as it only requires training an adapter for the new distortion and retraining the lightweight estimator.
 Prompt learning techniques, which capture task-specific context, have shown promise in guiding adaptation for vision tasks [24].
-Leveraging this concept, recent methods encode degradation-specific information and guide the restoration model to enhance low-fidelity images with more than one possible distortion.
-PromptIR [40], for instance, integrates a dedicated prompt block to capture degradation-specific features from input images, while DA-CLIP [30], MPerceiver [2], ProRes [31] and Painter [50] leverage large pre-trained models as prompt generators.
-The latter approaches, however, are often constrained by high memory demands of large models.
+In all-in-one IR, PromptIR [49] integrates a dedicated prompt block to capture degradation-specific features from input images, while DA-CLIP [37], MPerceiver [2], ProRes [39] and Painter [62] leverage large pretrained models as prompt generators.
+However, the latter approaches, as well as diffusion-based methods [81, 25, 35], are often constrained by the high memory demands of large models and iterative processes, respectively.
 Additionally, existing all-in-one approaches require all degradation types to be predefined during training, limiting the addition of new distortion to already trained models.
-In contrast, aiming at versatile image restoration, our model is developed to blindly process images with various kinds of distortion, either in single or composite forms.
-With the mixed design of lightweight degradation adapters, our model is also flexible and extensible to unseen distortions, thus shows great generalizability towards the challenging blind all-in-one IR problem.
+All-in-one IR also places great emphasis on adverse weather removal [60, 47, 84, 85], compression [77, 52, 33], video [80] or techniques to improve existing architectures [50, 66].
+In contrast, aiming at simple and versatile image restoration, we propose a three-phase approach to blindly process images with various degradations, either in single or composite forms.
+First, we introduce a synthetic pretraining stage to establish a strong baseline model.
+Next, we employ a specialize-then-merge strategy, where the model is first adapted to individual real-world degradations before learning to address the problem in a blind all-in-one manner.
 (a) Rain
@@
 (e) Low-light
-Figure 3: Examples of our synthetic degradation generation for five traditional distortions.
+(f) Deg. Cutmix
+Figure 3: Examples of our synthetic degradation generation for five traditional degradations (a)-(e) and our Degradation Cutmix (f).
 Parameter Efficient Fine-Tuning
-Fine-tuning models by updating all parameters is computationally inefficient, especially for large-scale models.
-Parameter Efficient Fine-Tuning (PEFT) addresses this by reducing the number of trainable parameters and memory usage while achieving comparable performance to full fine-tuning [54].
-Low-Rank Adaptation (LoRA) [19] introduced a reparametrization strategy for fine-tuned weights, where the weights of a specific layer $W^{\prime}=W+\Delta W$ are represented as a linear combination of the pre-trained weights $W$ and the updated term $\Delta W$ obtained through a low-rank decomposition, $\Delta W=BA$.
+Fine-tuning all model parameters is computationally expensive, especially for large-scale models, but Parameter Efficient Fine-Tuning (PEFT) addresses this by reducing the number of trainable parameters and memory usage while maintaining performance [67].
+Low-Rank Adaptation (LoRA) [23] introduced a reparametrization strategy for fine-tuning, where the weights of a specific layer $W^{\prime}=W+\Delta W$ are represented as a linear combination of the pretrained weights $W$ and the updated term $\Delta W$ obtained through a low-rank decomposition, $\Delta W=BA$.
 Here, $B\in\mathbb{R}^{d\times k}$, $A\in\mathbb{R}^{r\times k}$ with rank $r\ll\text{min}(d,k)$.
-By optimizing only the low-rank matrices $A$ and $B$, LoRA reduces the number of parameters required for adapting that layer (vs tuning all parameters), minimizing memory and computational demands.
-Building upon LoRA, alternative decomposition methods have been proposed, such as Vector-based Random Matrix Adaptation [24] and Conv-LoRA[67].
+By optimizing only the low-rank matrices $A$ and $B$, LoRA reduces the number of parameters required for adapting that layer, minimizing memory and computational demands.
+Building upon LoRA, alternative decomposition methods have been proposed, such as VeRA [29] and Conv-LoRA[82].
 PEFT techniques have proven valuable not only for domain adaptation in large models but also for applications like task arithmetic and continual learning, mitigating the problem of catastrophic forgetting — where models lose previously acquired knowledge upon learning new tasks [11].
@@
 [48] applied low-rank decomposition within single-task IR frameworks to enhance model performance.
-In contrast, we propose a simple but novel approach to merge different adaptors — LoRAs — to obtain a flexible blind all-in-one model capable of learning new tasks without forgetting previously learned IR tasks.
+Conversely, we propose a simple approach to merge different adaptors to obtain a flexible blind all-in-one model capable of learning new tasks without forgetting previously learned IR tasks.
 This is achieved through a disentangled scheme that requires minimal retraining for each new task.
@@
 Figure 2 presents an overview of our proposed approach for enhancing low-quality images by systematically addressing the key limitations in current IR methods.
-Our approach consists of three phases, roughly targeting the following challenges: (i) robust generalization to varying types of degradations (Phase I), (ii) effective adaptation to specific degradations (Phase II), (iii) flexible all-in-one mechanism able to handle unknown and mixed degradations (Phase III).
-Note that we freeze the parameters of baseline after Phase I and adapters after and II, thus preserving maximal knowledge learned from previous tasks and avoid catastrophic forgetting.
-This design also allows us to adapt to new distortions with minimal additional training as only a new adapter and the light-weighted estimator will be trained.
-Phase I: Pre-training with synthetic degradations.
-Recent advances demonstrate that large-scale pre-trained models can significantly improve performance across various tasks [1].
-Moreover, pre-training with synthetic data has also shown effectiveness when the domain gap is properly addressed [55].
-We hypothesize that applying large-scale pre-training to IR can yield a notable performance boost compared to training solely on traditional IR datasets.
-Thus, as depicted in Fig.
-2, Phase I involves training a baseline IR model with large-scale data with synthetic degradations to provide a robust weight initialization for handling complex degradations.
-To obtain large-scale data for pre-training, we define a degradation pipeline that generates a low-quality counterpart for each high-quality input image by introducing different synthetic distortions.
-We focus on five common distortions: noise, blur, rain, haze, and low-light conditions.
+Our approach consists of three phases, roughly targeting the following challenges: (i) robust generalization to varying types of degradations (Phase I), (ii) effective adaptation to specific degradations (Phase II), (iii) flexible blind all-in-one mechanism able to handle unknown and mixed degradations (Phase III).
+Note that we freeze the parameters of the baseline after Phase I and adapters after Phase II, preserving maximal knowledge learned from previous phases and avoiding catastrophic forgetting.
+This design also allows us to adapt to new distortions with minimal training as only a new adapter and the lightweight estimator will be trained.
+Phase I: pretraining with synthetic degradations.
+Recent advances demonstrate that large-scale pretrained models can significantly improve performance across various tasks [1].
+Moreover, pretraining with synthetic data has also shown effectiveness when the domain gap is properly addressed [68].
+We hypothesize that applying large-scale pretraining to IR can yield a notable performance boost compared to training solely on traditional IR datasets.
+Thus, as depicted in Figure 2, Phase I involves training a baseline IR model with large-scale data and synthetic degradations to provide a robust weight initialization.
+To obtain large-scale data for pretraining, we define a degradation pipeline that generates a low-quality counterpart for each high-quality input image by introducing different synthetic degradations.
+We pre-train on five distortions: noise, blur, rain, haze, and low-light conditions, as: (i) they are the most studied and commonly observed in real-world scenarios, (ii) they provide a diverse and balanced mix of low- and high-frequency degradations, and (iii) they enable a strong foundation for other distortions.
 Each synthetic distortion closely approximates its real-world counterpart.
@@
 In contrast, haze is introduced by adding an achromatic layer based on the depth map estimated from DepthAnythingv2 [68].
-Figure 3 illustrates examples of each synthetic degradation.
+Figure 3 (a)-(e) illustrate examples of each synthetic degradation.
 We use the Google Landmarks dataset (GLD) [65], a large-scale collection of five million images of diverse landmarks worldwide.
-To ensure that the inputs to our pipeline are high quality, we filter out the images with a resolution smaller than 400 pixels on any side and exclude images with a NIMA score [46] below the 4.90 threshold, retaining only high-quality natural images.
+To ensure that the input images are high quality, we filter out the images with a resolution smaller than 400 pixels and exclude images with a NIMA score [57] below the 4.90 threshold.
 To construct our baseline model, we build on recent advancements in all-in-one IR methods.
@@
 This combination enables the baseline model to effectively capture the inherent information of various degradations, establishing a strong foundation for generalized IR tasks.
-While training on single distortions enables the model to address each type individually, real-world IR often involves mixed degradations.
-However, directly applying multiple distortions to an image often leads to severe quality degradation, making it difficult for the model to learn to reconstruct the original image.
-To tackle this, we implement two strategies: (i) a degradation CutMix [61] technique, and (ii) a segmentation head with a cross-entropy loss.
-Our first strategy — inspired by Yang et al.
-[55] for depth estimation — applies two different degradations divided into different regions.
-This setup helps the model distinguish and manage multiple degradation types within the same image.
-As our second strategy, we incorporate a segmentation head into the baseline model, which outputs a per-pixel map of the distortions.
-By comparing this map with the ground truth, we guide the model to recognize and differentiate multiple types of degradation within a single image.
-We report additional details about our baseline and the generation of the synthetic distortions in the Supplementary Material.
-Table 1: 5-degradations setup.
-Quantitative results on five IR datasets comparing the state-of-the-art all-in-one methods and our approach.
-Ours (Oracle) is an upper bound for our approach: it computes the best reachable value in case our estimator always chooses the correct degradation.
+While training on single distortions allows the model to handle each type individually, real-world IR often involves mixed degradations.
+However, applying multiple synthetic degradations simultaneously can severely deteriorate image quality, making it harder for the model to reconstruct the original image (see Table 6).
+To address this, we implement two strategies: (i) a Degradation CutMix technique, and (ii) a segmentation head with a cross-entropy loss.
+In Figure 3 (f), we show our first strategy — inspired by Yang et al.
+[68] for depth estimation, where two different degradations are applied in separated random regions.
+Although this setup is not fully realistic, it helps the model distinguish and manage multiple degradation types within the same image.
+Our second strategy incorporates a segmentation head into the baseline model, producing a per-pixel distortion map.
+By comparing this map with the ground truth map, we guide the model to recognize and differentiate multiple types of degradation.
+The segmentation head is removed in subsequent phases.
+Additional details on our baseline and synthetic degradation generation are provided in the Supplementary Material.
+Table 1: 5-degradations setup. Quantitative results on five IR datasets comparing the state-of-the-art all-in-one methods and our approach.
 PSNR/SSIM Deraining Dehazing Denoising Deblurring Low-Light Average Param.
@@
 PromptIR [49] 35.40 .967 28.26 .965 30.89 .872 26.55 .808 21.80 .815 28.58 .885 36M
-Ours OH 37.73 .978 33.46 .983 31.38 .898 29.00 .878 24.20 .865 31.15 .920 59M
-Ours SW 37.79 .979 33.48 .984 31.38 .898 29.00 .878 24.19 .865 31.17 .921 59M
-Ours (Oracle) 39.09 0.981 33.54 .984 31.40 0.901 29.10 .879 24.45 .866 31.39 .922 59M
+AdaIR [10] 38.02 .981 30.53 .978 31.35 .889 28.12 .858 23.00 .845 30.20 .910 28M
+Ours OH 38.18 .983 33.46 .983 31.38 .898 29.00 .878 24.20 .865 31.24 .921 41M
+Ours SW 38.22 .984 33.48 .984 31.38 .898 29.00 .878 24.19 .865 31.25 .921 41M
 Phase II: Single-task adaptation.
-By generating distortions synthetically, we achieve a robust weight initialization for our baseline model.
-However, a domain gap persists between our pre-training data (due to both the different dataset and the synthetic distions), standard IR datasets, and real-world conditions.
-For instance, accurately simulating haze is particularly challenging due to its dependence on light scattering by particles, which varies with depth and atmospheric conditions.
-To bridge this gap, we propose adapting the baseline model with adapters learned for each specific task in its standard dataset, representing Phase II of our approach, as outlined in Fig.
-2.
-In our approach, each linear and convolutional layer is augmented with a set of LoRA, one per type of degradation, parametrized by $\left\{A_{n}\right\}$ and $\left\{B_{n}\right\}$ where $n$ is the index of the degradation.
-Following LoRA, these low-rank matrices adjust the initial frozen weights through linear combinations, as described earlier.
-For simplicity, we omit the subindex indicating the specific layer of the model, but each layer has a separate set of adapters.
-Due to the proposed pre-training approach and the subsequent specialization via adaptation, our approach achieves superior performance across diverse tasks and generalizes effectively.
+While synthetic degradations provide a robust weight initialization for our baseline model, a domain shift remains between the pretraining data (due to the different dataset and synthetic degradations), standard IR datasets, and real-world conditions.
+For instance, accurately simulating haze is particularly challenging, as it depends on light scattering by particles, which varies with depth and atmospheric conditions.
+To mitigate this gap, we introduce Phase II of our approach, where the baseline model is adapted using task-specific adapters specializing on a single degradation by using a standard IR dataset, as outlined in Figure 2.
+In this phase, we augment each linear and convolutional layer with a set of LoRA adapters, one per degradation type and parametrized by $\left\{A_{n}\right\}$ and $\left\{B_{n}\right\}$ where $n$ denotes the degradation index.
+These low-rank matrices refine the initial frozen weights through linear combinations and are trained individually for each degradation type.
+For simplicity, we omit the layer-specific indices, but each layer has its own set of adapters.
+By leveraging pretraining and subsequent specialization via adapters, our approach achieves superior performance across diverse tasks and exhibits improved generalization.
 Phase III: Multi-task integration.
-While LoRA adapters serve as a plug-and-play solution for pre-trained models, there is still the need to select the most suitable adapter based on the input image, when the type of degradation is unknown.
+LoRA adapters serve as a plug-and-play solution to refine our pretrained model.
+However, we still need to select the most suitable adapter based on the input image, when the degradation type is unknown.
 To address this limitation and derive a blind all-in-one IR method, we propose to use a lightweight degradation estimator $p\left(n|x;\theta\right)$ to estimate the probability of each degradation $n$ given the input image $x$, parametrized by $\theta$.
-This estimator is trained on the combination of all (seen) datasets, thus learning to identify the type of degradation present in the image.
-As shown in Fig.
-2, the probability of the estimator is used as weights to linearly combining the task-specific adapters with the baseline parameters.
+This estimator is trained to identify the type of degradation in the image using the combination of all the IR datasets in Phase II.
+As shown in Figure 2, the probabilities of the estimator are used to linearly combine the task-specific adapters with the baseline weights.
+In this way, the baseline is adapted with the task-specific LoRAs according to the lightweight estimator and, therefore, each input image individually.
 Specifically, given an input image $x$, the weight update for a specific layer with baseline weights $W$ and adapter weights $\left\{A_{n}\right\}$ and $\left\{B_{n}\right\}$ is computed as:
@@
 where $N$ is the number of seen degradations.
-We propose two variants of the estimator: (i) one-hot and (ii) soft-weights.
-The one-hot variant simply selects the adapter corresponding to the degradation with the maximum probability (equivalent to 1 with one-hot probabilities), while the soft-weight variant uses computes the weighted average as in Eq.
+We propose two variants of the estimator: (i) one-hot (OH) and (ii) soft-weights (SW).
+The one-hot variant simply selects the adapter corresponding to the degradation with the maximum probability, while the soft-weight variant computes the weighted average as in Eq.
 1.
-Our approach allows the model to linearly combine the degradation-specific knowledge to address both specific and multiple distortions.
-Its modular and parameter efficient architecture also allows to add additional degradations with small training effort: train additional adapters for the new task, and update the estimator.
+With a strong baseline model capable of generalizing effectively, our approach not only achieves superior performance on known degradations but also performs well on unseen and mixed degradations.
+Moreover, thanks to its modular architecture, we empirically demonstrate that performance on unseen degradation types can be further improved with minimal training effort.
+Specifically, one only needs to train an additional adapter for the new degradation type and retrain the lightweight estimator — together requiring fewer than four million parameters.
 Table 2: 3-degradations setup. Quantitative results on three IR datasets comparing the state-of-the-art all-in-one methods and our approach.
@@
 PromptIR [49] 36.37 .972 30.58 .974 33.98 .933 31.31 .888 28.06 .799 32.06 .913
-Ours OH 38.58 .981 33.71 .985 33.95 .934 31.29 .889 28.04 .798 33.11 .918
-Ours SW 38.52 .980 33.62 .984 33.95 .933 31.24 .889 28.01 .796 33.07 .916
+AdaIR [10] 38.64 .983 31.06 .980 34.12 .935 31.45 .892 28.19 .802 32.69 .918
+Ours OH 38.69 .982 33.53 .984 34.18 .935 31.38 .890 28.25 .804 33.21 .919
+Ours SW 38.64 .982 33.50 .984 34.18 .935 31.38 .890 28.22 .803 33.18 .919
 Table 3: Quantitative results on additional test datasets with the learned degradations.
@@
 DiffUIR [81] 14.78 .487 23.98 .739 26.12 .861
-Restormer [63] 14.50 .464 24.42 .781 27.12 .877
-PromptIR [40] 14.28 .444 24.49 .762 27.70 .870
-Ours OH 21.69 .692 27.04 .850 28.09 .907
-Ours SW 19.37 .594 27.05 .850 28.09 .906
+Restormer [76] 14.50 .464 24.42 .781 26.12 .877
+AdaIR [10] 14.13 .438 24.43 .771 26.54 .867
+PromptIR [49] 14.28 .444 24.49 .762 26.70 .870
+Ours OH 18.76 .612 27.04 .850 28.09 .907
+Ours SW 17.13 .564 27.05 .850 28.09 .906
 4 Experiments
-We evaluate our method on two setups for all-in-one IR: five- and three-degradation setups.
-Additionally, we test on datasets excluded from training, novel IR degradation types, and mixed degradation scenarios.
-The accuracy of the methods is assessed using two well-established metrics: PSNR and SSIM.
-In all cases we report the mean value for all the test images, and we highlight best and second-best values for each metric.
+We evaluate our method on two all-in-one IR setups: five- and three-degradation.
+Additionally, we assess its performance on datasets excluded from training, unseen IR degradation types, and mixed degradation scenarios.
+The accuracy of the methods is measured using two well-established metrics: PSNR and SSIM.
+In all cases we report the mean value across all the test images, and we highlight best and second-best values for each metric.
 Our approach is compared against recent all-in-one IR methods.
-Notably, some current methods are either evaluated under different setups or lack available code and models, which complicates direct comparisons.
-Therefore, we train three state-of-the-art methods, Restormer [63], PromptIR [40], and X-Restormer [8], on the five-degradation IR setup to provide a fair comparison.
-We used the authors’ provided code.
-For our approach, we evaluate the two variations based on the variants of the estimator, namely one-hot (OH) and soft weights (SW).
+However, some existing methods have been evaluated under different setups or lack publicly available training code and models [25, 2, 62].
+To ensure a fair comparison, we train five state-of-the-art methods, Restormer [76], DA-CLIP [37], DiffUIR [38], PromptIR [49], X-Restormer [8], on the five-degradation IR setup.
+We used the official code provided by the authors.
+For our approach, we evaluate the two variations, one-hot (OH) and soft weights (SW).
 Implementation details:
-Our training is conducted in three phases using Adam [23] with weight decay [29], an initial learning rate of $2\times 10^{-4}$, a cosine learning rate scheduler, and a warmup start of one epoch.
-In phase I, we apply a weighted combination of cross-entropy loss $L_{CE}$ for per-pixel degradation maps and a reconstruction loss $L_{R}$, where $L_{R}$ consists of $L_{1}$ loss and SSIM loss.
-Both the cross-entropy and SSIM components are weighted with 0.5 In phase II, we use only the reconstruction loss, where, as before, SSIM loss is weighted with 0.5.
-In phase III, we apply cross-entropy loss between the regressor’s output and the input image’s degradation type.
+Our training follows a three-phase scheme, using the Adam [28] optimizer with weight decay [36], an initial learning rate of $2\times 10^{-4}$, a cosine learning rate scheduler, and a warmup start of one epoch.
+Following Potlapalli et al.
+[49], we train our model on four NVIDIA A40 GPUs with a batch size of 4 and image patches of 384.
+In the Phase I, we jointly optimize a weighted combination of two objectives: a cross-entropy loss $L_{CE}$ for per-pixel degradation classification and a reconstruction loss $L_{R}$, which consists of $L_{1}$ loss and SSIM loss.
+The cross-entropy and SSIM components are both weighted at 0.5 to balance classification and reconstruction.
+During Phase II, we optimize only the reconstruction loss $L_{R}$, maintaining the same SSIM weighting as in the previous phase.
+In Phase III, we train the lightweight estimator separately using a cross-entropy loss to classify the degradation type of the input image, ensuring that the model accurately learns degradation prediction.
 Table 4: Quantitative results for unseen IR tasks.
@@
 Ours* shows results for the lightweight re-training scenario.
-New adapters are trained for the new tasks and the estimator is retrained with 8 tasks (5-IR case + 3 new ones; only 8M training parameters).
+New adapters are trained for the new tasks and the estimator is retrained with 8 tasks.
 PSNR/SSIM 4-to-8 bits JPEG Q20 Desnowing
@@
 Restormer [76] 24.64 .743 26.90 .929 18.14 .655
+AdaIR [10] 24.63 .739 26.65 .924 18.06 .649
 PromptIR [49] 24.70 .740 26.60 .920 18.49 .673
@@
 Ours SW 29.03 .810 30.71 .939 24.02 .779
-5-Degradation blind IR:
-For the blind five-task setup, we follow the protocol by Zhang et al.
+5-Degradation blind IR.
+For the blind five-task setup, we follow the protocol of Zhang et al.
 [78].
-Specifically, we use Rain200L[56] for deraining, RESIDE [25] for dehazing, BSD400 [33] and WED [32] for denoising with $\sigma=25$, GoPro [35] for deblurring, and LOL [52] for low-light enhancement.
-For evaluation, we employ Rain100L [56], SOTS-Outdoor [25], BSD68 [33], GoPro [35], and LOL [52].
-Results are reported in Tab.
-1.
-Our approach outperforms all state-of-the-art methods across all tasks, with the exception of PSNR for denoising.
-Notably, our method achieves a substantial improvement over the state of the art, yielding a 2.91 dB average PSNR gain across the five tasks.
-In particular, our method surpasses IDR [64] by 2.16 dB on image deraining and provides a 2.40 dB improvement over PromptIR [40] on low-light image enhancement.
-GoPro [35] Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-LoLv1 [52] Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Rain100H [56] Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Input Restormer [63] PromptIR [40] Ours-OH Ground truth
+Specifically, for training we use Rain200L[69] for deraining, RESIDE [30] for dehazing, BSD400 [41] and WED [40] for denoising with $\sigma{=}25$, GoPro [43] for deblurring, and LOL [64] for low-light enhancement.
+Evaluation is conducted on Rain100L [69], SOTS-Outdoor [30], BSD68 [41], GoPro [43], and LOL [64], with results reported in Table 1.
+We outperform all the methods across all tasks, except PSNR in denoising.
+Note that denoising gains are typically small due to the identical noise distribution in training and test sets in all-in-one setups.
+Our method achieves a substantial improvement over the state of the art average, with a PSNR gain of 1.05 dB.
+Remarkably, our method outperforms AdaIR [10] by 2.95 dB PSNR and 1.19 dB PSNR on dehazing and low-light image enhancement, respectively.
+GoPro [43] Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+LoLv1 [64] Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Rain100H [69] Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Input Restormer [76] PromptIR [49] AdaIR [10] Ours-OH Ground truth
 Figure 4: Qualitative results for single degradation removal, including deblurring on the GoPro [43] dataset, denoising on the LoLv1 [64] dataset, and deraining on the Rain100H [69] dataset.
-3-Degradation Blind IR:
+3-Degradation blind IR.
 Following Li et al.
-[27], we further evaluate our approach on a 3-task blind IR setup, comparing our approach with specialized all-in-one methods for deraining, dehazing, and denoising.
-Compared to the five-task setup, we omit deblurring and low-light enhancement and instead introduce two additional noise levels: $\sigma=15$ and $\sigma=50$.
-Two LoRA adapters for these noise levels are trained, and the estimator is retrained for the new setting.
-Results for this 3-task setup are presented in Tab.
-2.
-Our approach outperforms all the other state-of-the-art methods on average and stays consistent in all the degradation types.
-JPEG [42] Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-4-to-8 bits [42] Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Input PromptIR [40] Ours-SW Ours-SW retrained Ground truth
+[32], we evaluate our approach on a three-task blind IR setup, comparing it to specialized all-in-one methods for deraining, dehazing, and denoising.
+Compared to the five-task setup, we omit deblurring and low-light enhancement while introducing two additional noise levels: $\sigma{=}15$ and $\sigma{=}50$ on the BSD68 [41] dataset.
+Two LoRA adapters and the lightweight estimator are trained for these new noise levels.
+The baseline model remains the same, trained on the five-task setup.
+Results in Table 2 demonstrate that our approach outperforms all state-of-the-art methods on average while maintaining consistent performance across all degradation types.
+In detail, we achieve 0.52 dB average PSNR over AdaIR [10], even when our baseline was trained on the five-degradation setup.
+JPEG [53] Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+4-to-8 bits [53] Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Input PromptIR [49] AdaIR [10] Ours-SW Ours-SW retrained Ground truth
 Figure 5: Qualitative results for unseen IR tasks, including JPEG artifact removal and 4-to-8 bit reconstruction.
-PromptIR [40] and Ours are not trained for this task, while Ours retrained has a specified LoRA in an 8-degradation setup.
-Additional test sets:
-To assess the generalization capabilities of our approach, we evaluate the model on three datasets not used during training.
-Specifically, we use the Rain100H [56] dataset for heavy deraining, HIDE [43] for human-centric deblurring, and LoLv2-Real [58] for low-light image enhancement.
-The results of the three best methods on the 5-degradation setup and our approach are presented in Tab.
-3.
-Our method outperforms all other approaches across all datasets and metrics.
-Notably, we achieve a significant improvement of 7.19 dB over Restormer [63] on Rain100H [56].
-This performance boost can be attributed to the pre-training on synthetic data, which allows our model to learn from a broader set of natural images and degradation scenarios compared to the other methods’ small training set, like Rain200L [56].
-Additionally, we achieve over 1 dB PSNR improvement on both HIDE and LoLv2-Real, further demonstrating the effectiveness and robustness of our approach.
-Unseen IR tasks:
-Although our approach was initially trained on five types of degradations, we further evaluate its generalization to three additional IR tasks not encountered during training.
-Specifically, we assess our method on JPEG artifact removal and bit-depth reconstruction using the Live1 dataset [42], and on desnowing using CityScapes-Snow-Medium [65].
-Results for the top four previous state-of-the-art methods from the five-degradation setup, alongside our approach, are presented in Tab.
-4.
-Remarkably, our method outperforms all other approaches across all tasks and metrics.
-To show the adaptability of our approach, we additionally train a separate adapter for each new task, along with a new estimator that blends all eight task-specific adapters, thereby expanding the model’s versatility.
-The last two rows of Tab.
-4 report results for the newly integrated tasks — shown with a .
-Notably, with the addition of these task-specific adapters, our model achieves a notable boost in performance, while maintaining an average PSNR of 30.08 dB across the original five IR tasks, outperforming the other state-of-the-art methods trained on the initial setup.
-Blur&Noise Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Haze&Snow Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Input Restormer [63] PromptIR [40] Ours-SW Ground truth
+PromptIR [49], AdaIR [10], and Ours are not trained for these tasks, while Ours-retrained has a specified LoRA in an 8-degradation setup.
+Additional test sets.
+To assess generalization, we evaluate our model on three datasets not used during training: Rain100H [69] for heavy deraining, HIDE [54] for human-centric deblurring, and LoLv2-Real [71] for low-light image enhancement.
+Table 3 presents the results for the methods trained in the five-task setup.
+Our method outperforms all competitors across all datasets and metrics, achieving substantial improvements of 3.98 dB over DiffUIR [81] on Rain100H, 1.65 dB over X-Restormer [8] on HIDE, and 1.39 dB over PromptIR [49] on LoLv2-Real.
+These results highlight the effectiveness and robustness of our approach, which benefits from pretraining on synthetic degradations, enabling it to generalize across diverse natural images and degradation scenarios.
+Unseen IR tasks.
+Although our model is trained on five degradation types, we assess its adaptability to three additional IR tasks: JPEG artifact removal and bit-depth reconstruction using the Live1 dataset [53], and desnowing using CityScapes-Snow-Medium [79].
+The results in Table 4 demonstrate that our method consistently outperforms competing approaches.
+Moreover, by leveraging its modular scheme, we train independent LoRA adapters for each new task, blending them with the five-task baseline (denoted with in Table 4).
+This strategy allows for refining the model with minimal training (3.6M parameters per task) while preserving previously acquired knowledge, enabled by our robust baseline model and three-phase training procedure.
+In contrast, other methods either require retraining the entire model from scratch or fine-tuning it without ensuring that prior knowledge is retained.
+Blur & N Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Haze & Snow Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Input Restormer [76] PromptIR [49] AdaIR [10] Ours-SW Ground truth
 Figure 6: Qualitative results for two examples of mixed degradations.
@@
 Columns display the input image, results from Restormer [76], PromptIR [49], our method, and the ground truth, respectively.
-Mixed degradations:
-The limited availability of datasets with mixed degradations presents a challenge for training IR models under these complex conditions.
-To evaluate our approach, we evaluate it on three mixed-degradation scenarios: blur combined with noise $\sigma{=}25$ using the GoPro dataset[35], blur with JPEG artifacts using the REDS dataset [36], and haze combined with snow using the SRRS dataset [6].
-Notably, some degradations, such as JPEG artifacts and snow, are unseen during training, further increasing the difficulty of these tests.
-As shown in Tab.
-5, our model consistently outperforms other methods across all scenarios and metrics.
-Handling blur with noise is particularly challenging due to their contrasting nature: blur affects low frequencies, while noise affects high frequencies.
-Even though, our model achieved around 2dB improvement while others remained at the level of 22dB.
+Mixed degradations.
+We also evaluate our approach on three mixed-degradation scenarios: blur with noise ($\sigma{=}25$) using GoPro [43], blur with JPEG artifacts using REDS [44], and haze with snow using SRRS [5].
+Notably, JPEG artifacts and snow are unseen during training.
+As shown in Table 5, our model outperforms all competitors by at least 1 dB PSNR in all three cases.
+Addressing blur and noise simultaneously is particularly challenging due to their opposing frequency characteristics.
+Our specialize-then-merge scheme allows the lightweight estimator to efficiently blend the task-specific adaptors, offering greater flexibility than end-to-end models trained on fixed degradation types.
+Qualitative results.
+In Figure 4, we present examples featuring three types of single degradations from the five-task setup.
+Our approach effectively enhances the license plate in the first row, reduces noise in low-light conditions in the second row, and removes heavy rain streaks in the third.
+In Figure 5, we show unseen IR tasks including JPEG artifact removal and 4-to-8-bit reconstruction using the Live1 dataset [53].
+Our method successfully removes JPEG artifacts from the parrot’s plumage and artifacts caused by bit-depth reduction from the cloud and blue sky.
+By training a new adapter for these tasks (ours-SW retrained), we achieve superior results with minimal additional training.
+Finally, in Figure 6, we demonstrate our method’s performance under mixed degradations.
+The first row shows that our method achieves the best reconstruction of the text in the image, while the second row illustrates our model’s effectiveness in removing haze and snowflakes.
 Table 5: Quantitative results on datasets with mixed degradations.
@@
 Restormer [76] 22.35 .662 23.24 .698 20.76 .800
+AdaIR [10] 22.45 .663 23.16 .689 20.77 .802
 PromptIR [49] 22.89 .671 23.92 .705 20.94 .803
-X-Restormer [8] 22.67 .669 23.98 .710 20.76 .805
 Ours OH 24.30 .743 24.81 .717 21.48 .834
 Ours SW 25.14 .750 24.97 .719 22.09 .839
-Table 6: Ablation studies on types of pre-training for Phase I, and the rank of LoRA [19] for Phase II and III.
-Pre-training PSNR SSIM
-IR datasets 28.50 .892
-GLD+synth. 30.63 .913
-+ CutMix 31.09 .920
-+ Aux. segm. 31.17 .921
-Rank PSNR SSIM Params
-4 31.17 .921 3.6M
-8 31.14 .920 7.2M
-16 30.97 .916 14.3M
-Qualitative results:
-In Fig.
-4, we present examples featuring three types of single degradations from the five-task setup on which the model was trained.
-Our approach effectively enhances the license plate in the first row, reduces noise in low-light conditions in the second row, and removes heavy rain streaks in the third.
-In Fig.
-5, we show unseen IR tasks including JPEG artifact removal and 4-to-8-bit reconstruction using the Live1 dataset [42].
-Our method successfully removes JPEG artifacts from the parrot’s plumage and artifacts caused by bit-depth reduction from the cloud and blue sky.
-By training a new adapter for these tasks (ours-SW retrained), we achieve superior results with minimal additional training time.
-Finally, in Fig.
-6, we demonstrate our method’s performance under mixed degradations.
-The first row shows that our method achieves the best reconstruction of the text in the image, while the second row illustrates our model’s effectiveness in removing both haze and snowflakes.
-Ablation studies:
-In Tab.
-6, we perform an ablation study to assess the contribution to the final performance of the proposed pre-training pipeline and the LoRAs’ rank.
-For pre-training in Phase I, we use the same setup as the five-task configuration, i.e.
-training the LoRAs and the estimator using task-specific IR datasets and the obtained pre-trained weights.
-The first row — IR datasets — reflects pre-training on standard IR benchmarks, yielding performance comparable to state-of-the-art methods — see Tab.
-1.
-In contrast, simply pre-training on GLD with synthetic degradations improves performance by 2.15 dB PSNR, demonstrating the importance of Phase I to provide the model with strong generalization capabilities.
-Further incorporating both the CutMix-like [61] strategy and the auxiliar segmentation task provides an additional performance boost.
-We also compare LoRA [19] with different ranks for Phase II and III.
-The model achieves the best performance and efficiency when the rank is set to 4, which we adopted as our default setting.
-For more ablation results, please refer to our Supplementary Material.
+Table 6: Ablation studies on pretraining strategies for our method (a), and PromptIR [49] (b). We report the five-task setup average.
+(a) ABAIR PSNR SSIM
+IR datasets 29.50 .892
+GLD+2 global 27.68 .887
+GLD+1 global 30.63 .913
+GLD+CutMix 31.11 .920
++ Aux. seg. 31.21 .921
+(b) PSNR SSIM
+PromptIR [49] 28.58 .885
+RAM [50] 28.79 .889
+Art [66] 28.83 .889
+GLD+1 global 29.27 .902
+GLD+CutMix 29.58 .908
+Ablation studies.
+Our method consistently outperforms previous state-of-the-art approaches across seen, unseen, and mixed degradations, as demonstrated in all the tables.
+To further assess its effectiveness, we conduct additional experiments regarding the impact of our Phase I pretraining strategy.
+Table 6 reports the results averaged over the five-task setup.
+First, we evaluate the impact of the synthetic pretraining to the final performance of our method.
+Table 6 (a) compares different pretraining setups: standard IR datasets as pretraining (IR datasets), Google Landmarks dataset with one and two simultaneous global synthetic degradations applied across the entire image, (GLD+1 global and GLD+2 global), respectively, as well as incorporating the Degradation CutMix (GLD+Cutmix) and the auxiliary segmentation head with its corresponding loss (Aux.
+seg.).
+Next, we compare our synthetic pretraining approach with the pretraining strategy from RAM [50] and the optimization technique from Art [66] on the PromptIR [49] model.
+We select PromptIR as a well-established baseline for evaluation.
+The results of these comparisons are presented in Table 6 (b).
+From Table 6, we derive four key observations: (i) Our synthetic pretraining substantially outperforms other pretraining and optimization techniques.
+(ii) Our three-phase approach is effective, as even without using synthetic pretraining (IR datasets), our method surpasses the baseline PromptIR model.
+(iii) Training with multiple simultaneous degradations (GLD+2 global) negatively impacts performance, likely due to excessive degradation in Phase I, which hinders the learning of robust baseline.
+(iv) Both Degradation CutMix and the auxiliary segmentation head contribute substantially to performance improvements.
+These findings highlight the importance of both synthetic pretraining and the three-phase training scheme in achieving optimal results.
+Additional results and ablation studies for Phases II and III, including analyses of different adapter types, ranks, and alternative strategies for blending the adapters, are provided in the Supplementary Material.
 5 Conclusion
-In this work, we introduced an adaptive blind all-in-one IR model aiming towards practical IR.
-We design both dedicated adapters for strongly handling specific distortions but also a flexible architecture for dealing with practical IR challenges.We first developed a pre-training pipeline featuring multiple synthetic degradations over a large dataset that boosts the generalization of the model.
+In this work, we introduce an adaptive blind all-in-one IR (ABAIR) model aiming towards practical IR.
+We design a specialize-then-merge scheme with dedicated adapters for strongly handling specific distortions and a flexible architecture for dealing with unseen and mixed degradations.
+We first developed a pretraining pipeline featuring multiple synthetic degradations over a large dataset that boosts the model’s generalization.
 Second, we derived compact per-task adapters that robustly adapt to specific degradations.
-Third, we developed a lightweight degradation estimator that identifies varying degradations to blend the respective adapters, which is also capable of efficiently incorporate new degradations by training a small fraction of parameters.
+Third, we developed a lightweight degradation estimator that identifies varying degradations to blend the respective adapters.
+Our method is also capable of efficiently incorporate new degradations by training a small fraction of parameters.
 Our model largely outperforms the state-of-the-art on five- and three-task IR setups, and shows improved generalization to unseen datasets and IR tasks.
@@
 As with rainy images, capturing paired degraded and clean images under hazy conditions is nearly impossible due to the variability of environmental factors.
-Inspired by prior works [26, 3, 16], we model haze degradation using the Kochsmieder model [17], which describes how the visibility of distant objects diminishes, vanishing into the horizon as a function of their distance from the observer.
+Inspired by prior works [31, 17], we model haze degradation using the Kochsmieder model [21], which describes how the visibility of distant objects diminishes, vanishing into the horizon as a function of their distance from the observer.
 We can formulate it as:
 $I_{haze}=T\cdot I_{input}+(1-T)\cdot A,$ (3)
-where $I_{input}$ is the input clean image, $T$ is the transmission map derived from the estimated depth map, $A$ is the airlight, and $I_{haze}$ is the resulting image with synthetic haze.
+where $I_{input}$ is the input clean image, $T$ is the transmission map derived from the estimated depth map, $A$ is the haze color, and $I_{haze}$ is the resulting image with synthetic haze.
 We estimate the depth map of the input image using DepthAnythingv2 [68].
-This depth map is normalized within the predefined minimum and maximum haze values to generate the transmission map $T$.
+This transmission map is normalized within the predefined minimum and maximum haze values to generate the transmission map $T$.
 The depth map is then replicated for each color channel to compute Eq.
@@
 To ensure variability, all parameters are randomly sampled for each forward pass.
-The parameter ranges are as follows: minimum haze $[0.2,0.4]$, and maximum haze $[0.7,0.9]$, airlight is unique for the three color channels with values ranging $[140,200]$.
+The parameter ranges are as follows: minimum haze $[0.2,0.4]$, and maximum haze $[0.7,0.9]$, haze color is unique for the three color channels with values ranging $[140,200]$.
 Blur:
@@
 $I_{noise}=I_{input}+\mathcal{N}(0,\sigma).$ (5)
-Low-Light:
-When an image is captured under low-light conditions, cameras amplify the sensor’s signal to capture details, amplifying noise, and leading to grainy or speckled artifacts.
-Furthermore, the reduced dynamic range limits the ability to capture the range of intensities, resulting in color inconsistencies, and loss of detail in shadows and highlights.
-To simulate these conditions, we compress the input image histogram by a factor $c$ and add a noise with a meager $\sigma$ value.
-This process can be expressed mathematically as:
-$I_{lol}=I_{input}\cdot c+\mathcal{N}(0,\sigma).$ (6)
-The compression factor $c$ ranges in the interval $[0.25,0.5]$ and the $\sigma$ takes values in the interval $[0.5,1.5]$.
 Table 7: Ablation study on different low-rank adapters and their rank. Results are the mean for all images. LoRA outperforms both VeRA and Conv-LORA. Lower ranks perform better.
@@
 8 36.94 .968 32.44 .968 31.30 .895 28.48 .868 23.62 .575 30.56 .855 7.5M
+Table 8: Ablation study on different methods for blending the five degradations task-specific LoRA [23] adapters.
+PSNR/SSIM Deraining Dehazing Denoising Deblurring Low-Light Average
+Method Rain100L SOTS (Out) BSD68 σ=25 GoPro LoLv1
+Sum 19.50 .755 18.80 .745 18.10 .730 18.25 .728 17.70 .715 18.67 .736
+Average 30.54 .939 20.87 .855 28.98 .785 21.34 .792 15.49 .673 23.84 .809
+Ours OH 38.18 .983 33.46 .983 31.38 .898 29.00 .878 24.20 .865 31.24 .921
+Ours SW 38.22 .984 33.48 .984 31.38 .898 29.00 .878 24.19 .865 31.25 .922
+Ours Oracle 39.09 .981 33.54 .984 31.40 .901 29.10 .879 24.45 .866 31.39 .913
+Low-Light:
+When an image is captured under low-light conditions, cameras amplify the sensor’s signal to capture details, amplifying noise, and leading to grainy or speckled artifacts.
+Furthermore, the reduced dynamic range limits the ability to capture the range of intensities, resulting in color inconsistencies, and loss of detail in shadows and highlights.
+To simulate these conditions, we compress the input image histogram by a factor $c$ and add a noise with a small $\sigma$ value.
+This process can be expressed mathematically as:
+$I_{lol}=I_{input}\cdot c+\mathcal{N}(0,\sigma).$ (6)
+The compression factor $c$ ranges in the interval $[0.25,0.5]$ and the $\sigma$ takes values in the interval $[0.5,1.5]$.
 Appendix C Additional Phase II Ablation Studies
@@
 Regarding the decomposition rank, we find that ranks of 4 and 8 consistently outperform a rank of 16 in terms of both accuracy and parameter efficiency.
-Table 8: Ablation study on different methods for blending the five degradations task-specific LoRA [19] adapters.
-PSNR/SSIM Deraining Dehazing Denoising Deblurring Low-Light Average
-Method Rain100L SOTS (Out) BSD68 σ=25 GoPro LoLv1
-Sum 19.50 .755 18.80 .745 18.10 .730 18.25 .728 17.70 .715 18.67 .736
-Average 30.54 .939 20.87 .855 28.98 .785 21.34 .792 15.49 .673 23.84 .809
-Ours OH 37.73 .978 33.46 .983 31.38 .898 29.00 .878 24.20 .865 31.15 .920
-Ours SW 37.79 .979 33.48 .984 31.38 .898 29.00 .878 24.19 .865 31.17 .921
 Appendix D Additional Phase III Analysis
-In the third phase of our method, we train an estimator to select or blend the most suitable set of adapters based on the input image.
+In the third phase of our method, we train an estimator to integrate the set of adapters with the baseline model based on the input image.
 The estimator architecture comprises four blocks of Conv2D layers, each followed by batch normalization, ReLU activation, and max pooling, culminating with a global average pooling layer and a linear projection.
@@
 Notably, even when the estimator selects an incorrect adapter, no significant artifacts or undesired effects are introduced, owing to the robustness of the large-scale pre-training.
-Finally, to highlight the importance of our estimator, we perform a last ablation comparing our approach to just adding or averaging the five degradation task-specific adapters.
+To highlight the importance of our estimator, we perform another ablation comparing our approach to just adding or averaging the five degradation task-specific adapters.
 Tab.
@@
 On the other hand, averaging the adapters results in suboptimal restored images, as the model fails to specialize in any particular degradation type.
+Finally, Table 8 also presents the results for the five-task setup assuming the lightweight estimator is perfect, i.e., an oracle.
+The results are only slightly better from those obtained when using our estimator, confirming that its performance is strong.
+Please note that these Oracle results can be interpreted as single-task performance, as the selected adapter always corresponds to the specific degradation type of the input image.
 Figure 10: Confusion matrix for our estimator in test images.